# Chen, Yan, and Jiang

### v1

@@ 4 Application on Pricing
In this section, we leverage the (weak) concavity of $c(\boldsymbol{\lambda})$ to design an efficient algorithm for solving the upper-level pricing problem. We assume a one-to-one mapping between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$, as well as concavity of the revenue function with respect to the arrival rate. Both assumptions are standard in the revenue management literature. {assumption} There exists a one-to-one correspondence between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$. The revenue function $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
@@ 4.2 Chicago Ridesharing Case Study
In this subsection, we conduct numerical experiments using the Chicago ridesharing dataset. The dataset, provided by Chicago Data Portal, [^1] https://data.cityofchicago.org/Transportation/Transportation-Network-Providers-Trips-2018-2022-/m6dm-c72p/about_data contains trips reported by ridesharing companies within the city of Chicago. The dataset divides the city (excluding the O’Hare International Airport) into 76 community areas, and each data point includes the passenger’s OD pair, represented by the corresponding communities’ centroid locations. Same as Yan, Yan, and Shen (2025), we focus on the typical rush-hour data (i.e., 7:30 AM - 8:30 AM on each Monday) in October and November 2019, before the COVID-19 pandemic, to conduct the experiment. In total seven weeks are selected as offline data, for the purpose of solving program (1) and obtaining the optimal static price.
To aggregate riders into distinct types, we apply the $K$-means clustering algorithm to the origin-destination (OD) pairs of all trips, with the number of clusters set to $N\in\{100,150,200\}$. For each cluster (i.e., rider type $i\in[N]$), we treat the cluster center as the representative OD for that type, and the average number of requests per hour within the cluster as the maximum possible hourly arrival rate $\overline{\lambda}_{i}$. Given the OD of each rider type, we compute the corresponding solo trip distances and pooled trip distances (in miles), denoted as $[l_{(i)}]_{i\in[N]}$ and $[l_{(i,j)}]_{i,j\in[N]}$, respectively. We let $c_{(i,j)}=cl_{(i,j)}$, where $c$ is the per-mile cost. We consider three possible values of $c\in\{0.7,0.9,1.1\}$ in the experiment, with a unit of dollar per mile. We assume in this case study that the riders’ per-mile willingness to pay is uniformly distributed in $[0,1]$, then we have $p_{i}(\lambda_{i})=l_{(i)}\left(1-\lambda_{i}/\overline{\lambda}_{i}\right)$, $\forall i\in[N]$.
@@
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 908 921 634 637 687 692 699 740 699 834 794 810
PG ($\delta_{\text{PG}}=10$) 332 415 375 374 304 321 298 330 704 863 657 633
PG ($\delta_{\text{PG}}=100$) 396 366 391 416 346 349 354 350 806 867 800 660
MM 64 68 59 63 63 71 75 79 124 158 194 171
Num. Iterations PG ($\delta_{\text{PG}}=1$) 70.3 72.3 47.3 48.3 57.7 59.7 60.3 57.0 56.7 70.7 67.3 68.0
PG ($\delta_{\text{PG}}=10$) 24.7 32.7 27.7 28.0 25.0 26.3 25.3 27.0 54.0 68.7 54.7 54.0
PG ($\delta_{\text{PG}}=100$) 29.3 28.7 30.0 30.3 28.3 29.0 30.0 30.0 59.0 65.3 67.0 55.3
MM 5.3 5.7 5.0 5.3 5.3 6.0 6.3 6.7 10.3 13.3 16.3 14.7
Objective Function PG ($\delta_{\text{PG}}=1$) 64.13 58.92 45.75 37.96 39.45 33.70 20.65 13.88 16.68 13.05 4.69 1.63
PG ($\delta_{\text{PG}}=10$) 64.33 59.11 45.93 38.14 39.64 33.89 20.85 14.11 13.11 10.21 4.31 1.64
PG ($\delta_{\text{PG}}=100$) 64.33 59.11 45.93 38.14 39.64 33.89 20.85 14.11 12.58 10.08 4.43 1.64
MM 64.33 59.11 45.93 38.14 39.64 33.89 20.84 14.11 19.59 14.30 4.97 1.71
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 829 747 551 658 720 748 756 542 1200 829 919 735
PG ($\delta_{\text{PG}}=10$) 772 743 710 758 737 760 765 760 1036 959 1099 994
PG ($\delta_{\text{PG}}=100$) 1008 949 921 903 886 896 936 962 1200 1200 1024 1133
MM 132 133 133 134 159 179 170 162 411 402 446 292
Num. Iterations PG ($\delta_{\text{PG}}=1$) 30.7 27.7 20.7 22.3 25.0 27.7 28.7 20.0 44.3 30.3 33.7 25.3
PG ($\delta_{\text{PG}}=10$) 26.7 26.7 26.7 27.0 26.3 27.3 28.3 28.0 39.3 35.3 40.0 34.7
PG ($\delta_{\text{PG}}=100$) 33.7 33.3 33.3 33.3 32.0 31.7 34.3 35.0 44.7 43.7 38.0 39.7
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.3 6.0 15.7 15.3 17.0 11.0
Objective Function PG ($\delta_{\text{PG}}=1$) 61.55 56.25 44.15 37.47 36.35 30.71 19.45 13.64 15.36 11.74 4.56 1.65
PG ($\delta_{\text{PG}}=10$) 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 12.98 10.13 4.38 1.58
PG ($\delta_{\text{PG}}=100$) 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 13.60 10.67 4.22 1.52
MM 61.71 56.32 44.30 37.55 36.40 30.79 19.52 13.73 16.59 12.09 4.66 1.72
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1069 1076 1078 1112 1001 1069 1058 1123 1200 948 1104 1015
PG ($\delta_{\text{PG}}=10$) 1183 1185 1185 1206 1137 1200 1200 1200 1200 1202 1200 1200
@@
MM 238 254 254 242 302 301 286 293 704 640 563 502
Num. Iterations PG ($\delta_{\text{PG}}=1$) 22.3 22.3 22.3 22.3 20.7 22.0 21.3 21.3 24.0 19.3 22.7 20.3
PG ($\delta_{\text{PG}}=10$) 25.3 25.3 25.0 24.3 23.0 25.3 24.3 24.7 24.3 24.7 24.7 23.3
PG ($\delta_{\text{PG}}=100$) 24.7 24.7 24.7 25.0 25.0 25.0 24.7 24.7 24.7 24.0 24.3 24.7
MM 5.0 5.3 5.3 5.0 6.3 6.3 6.0 6.0 15.0 13.3 11.7 10.3
Objective Function PG ($\delta_{\text{PG}}=1$) 59.78 54.82 43.69 37.29 34.39 29.31 18.78 13.47 14.59 11.08 4.35 1.52
PG ($\delta_{\text{PG}}=10$) 59.97 54.88 43.75 37.35 34.36 29.34 18.94 13.45 13.80 10.64 3.97 1.53
PG ($\delta_{\text{PG}}=100$) 55.46 50.38 31.77 29.41 32.36 21.50 6.88 -5.98 1.35 -2.16 -2.99 -3.02
MM 59.97 54.89 43.77 37.37 34.48 29.37 19.06 13.54 15.23 11.34 4.48 1.61
Table 3 presents the results of the first set of experiments, where we assume $\theta_{1}=\cdots=\theta_{N}=\theta\in\{1/5,1/3,1,2\}$ per minute. Accordingly, the mean sojourn times for riders are $\{5,3,1,0.5\}$ minutes, respectively. The results demonstrate that the MM algorithm substantially reduces both the running time and the number of iterations compared to the PG method. Compared to PG with three different step sizes ($\delta_{\text{PG}}=1,10,100$), the MM algorithm reduces average running time by 72.6%, 71.6%, and 73.8%, respectively, and reduces the average number of iterations by 77.5%, 72.9%, and 75.2%. Since the running time is capped at 20 minutes, the actual time savings may be even greater. Moreover, within the maximum allowed running time, the MM algorithm consistently achieves a higher objective value—on average, 1.05%, 2.05%, and 14.27% higher than the PG methods. We also observe that the performance of PG is sensitive to step size selection: although a step size of 1 yields the best average performance, as shown in Table 4, it is not optimal in every instance. For example, a step size of 10 performs better when $c=0.7$ or $c=0.9$. In contrast, MM demonstrates strong robustness and outperforms PG across all cases and step-size rules.
@@
PG ($\delta_{\text{PG}}=1$) 50.14 25.31 8.41
PG ($\delta_{\text{PG}}=10$) 50.28 25.42 7.36
PG ($\delta_{\text{PG}}=100$) 47.87 21.97 4.33
MM 50.28 25.45 9.02
It can also be observed that as the number of types $N$ increases, the running time of both GA and MM increases, primarily because solving $c(\boldsymbol{\lambda})$ involves $N+N^{2}$ variables at each iteration. However, the number of iterations required by the MM algorithm does not scale proportionally with $N$, and in many cases remains nearly constant at a single-digit value. Even when $N=200$, the MM algorithm remains efficient and consistently produces results within a short CPU time.
@@
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 226 302 226 243 327 313
PG ($\delta_{\text{PG}}=10$) 352 358 321 314 737 406
PG ($\delta_{\text{PG}}=100$) 357 358 367 367 760 682
MM 60 59 72 84 152 134
Num. Iterations PG ($\delta_{\text{PG}}=1$) 19.0 24.7 19.0 20.3 26.7 26.3
PG ($\delta_{\text{PG}}=10$) 29.3 29.7 27.0 26.3 59.0 34.3
PG ($\delta_{\text{PG}}=100$) 30.0 30.3 30.7 30.7 61.0 56.7
MM 5.0 5.0 6.0 7.0 12.7 11.0
Objective Function PG ($\delta_{\text{PG}}=1$) 61.72 41.75 36.82 17.16 15.29 2.99
PG ($\delta_{\text{PG}}=10$) 61.73 41.76 36.82 17.17 12.92 2.83
PG ($\delta_{\text{PG}}=100$) 61.75 41.76 36.82 17.17 11.39 2.79
MM 61.75 41.76 36.82 17.17 17.00 3.04
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 981 576 594 758 1050 831
PG ($\delta_{\text{PG}}=10$) 734 734 742 757 1020 902
PG ($\delta_{\text{PG}}=100$) 921 898 841 897 1110 1102
MM 134 144 161 161 403 425
Num. Iterations PG ($\delta_{\text{PG}}=1$) 37.0 21.7 22.7 28.0 39.3 30.3
PG ($\delta_{\text{PG}}=10$) 27.3 27.7 27.3 27.7 38.0 32.7
PG ($\delta_{\text{PG}}=100$) 34.3 33.3 31.0 33.3 41.3 40.7
MM 5.0 5.3 6.0 6.0 14.3 15.0
Objective Function PG ($\delta_{\text{PG}}=1$) 58.71 40.30 33.25 16.12 13.23 2.72
PG ($\delta_{\text{PG}}=10$) 58.85 40.47 33.36 16.22 11.81 2.67
PG ($\delta_{\text{PG}}=100$) 58.85 40.47 33.36 16.22 11.79 2.65
MM 58.85 40.47 33.36 16.22 14.04 2.92
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1042 1086 1099 1100 1188 1200
PG ($\delta_{\text{PG}}=10$) 1210 1205 1200 1200 1200 1200
@@
MM 238 240 305 302 875 826
Num. Iterations PG ($\delta_{\text{PG}}=1$) 22.0 22.7 23.0 22.7 24.3 25.0
PG ($\delta_{\text{PG}}=10$) 25.7 25.3 25.0 25.7 25.0 25.0
PG ($\delta_{\text{PG}}=100$) 25.7 25.3 25.0 26.0 25.0 25.0
MM 5.0 5.0 6.3 6.0 18.3 17.0
Objective Function PG ($\delta_{\text{PG}}=1$) 56.90 39.78 31.31 15.48 12.15 2.25
PG ($\delta_{\text{PG}}=10$) 57.28 40.02 31.71 15.90 12.07 2.60
PG ($\delta_{\text{PG}}=100$) 53.44 33.93 26.24 14.68 5.87 -3.64
MM 57.28 40.14 31.73 15.91 13.05 2.78
Finally, Table 5 presents the experimental results for the setting where $\theta_{i}$ varies across agent types. For type-$i$ agents, $\theta_{i}$ is randomly drawn from a uniform distribution $U[\underline{\theta},\overline{\theta}]$, with two parameter choices: $(\underline{\theta},\overline{\theta})\in\{(1/5,1/3),(1,2)\}$. The results exhibit a similar trend to that observed in Table 1. Compared to three PG variants with $\delta_{\text{PG}}=1,10,100$, the MM algorithm reduces running time by 63.7%, 67.3% and 69.9%; decreases the number of iterations by 65.7%, 71.0% and 74.2%; and improves the objective value by 1.27%, 1.63% and 8.32% within the 20-minute maximum runtime, respectively. The performance of GA is also sensitive to the step size: $\delta_{\text{GA}}=10$ performs better when $c\leq 0.9$, whereas $\delta_{\text{GA}}=1$ is preferable when $c=1.1$.

### v2

@@ 4 Application on Pricing
In this section, we leverage the (weak) concavity of $c(\boldsymbol{\lambda})$ to design an efficient algorithm for solving the upper-level pricing problem. We assume a one-to-one mapping between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$, as well as concavity of the revenue function with respect to the arrival rate. Both assumptions are standard in the revenue management literature. {assumption} There exists a one-to-one correspondence between $\lambda_{i}$ and $p_{i}$ for all $i\in[N]$. The revenue function $\sum_{i\in[N]}\lambda_{i}p_{i}(\lambda_{i})$ is concave on $\Lambda$.
@@ 4.2 Chicago Ridesharing Case Study
In this subsection, we conduct numerical experiments using the Chicago ridesharing dataset. The dataset, provided by Chicago Data Portal, [^1] https://data.cityofchicago.org/Transportation/Transportation-Network-Providers-Trips-2018-2022-/m6dm-c72p/about_data contains trips reported by ridesharing companies within the city of Chicago. The dataset divides the city (excluding the O’Hare International Airport) into 76 community areas, and each data point includes the passenger’s OD pair, represented by the corresponding communities’ centroid locations. We use typical rush-hour data (i.e., 7:30 AM - 8:30 AM on Mondays), spanning 7 weeks between October 7 and November 18, 2019, prior to the COVID-19 pandemic, to conduct the experiment.
To aggregate riders into distinct types, we apply the $K$-means clustering algorithm to the origin-destination (OD) pairs of all trips, with the number of clusters set to $N\in\{100,150,200\}$. For each cluster (i.e., rider type $i\in[N]$), we treat the cluster center as the representative OD for that type, and the average number of requests per hour within the cluster as the maximum possible hourly arrival rate $\overline{\lambda}_{i}$. Given the OD of each rider type, we compute the corresponding solo trip distances and pooled trip distances (in miles), denoted as $[l_{(i)}]_{i\in[N]}$ and $[l_{(i,j)}]_{i,j\in[N]}$, respectively. We let $c_{(i,j)}=cl_{(i,j)}$, where $c$ is the per-mile cost. We consider three possible values of $c\in\{0.7,0.9,1.1\}$ in the experiment, with a unit of dollar per mile. We assume in this section that the riders’ per-mile willingness to pay is uniformly distributed in $[0,1]$, then we have $p_{i}(\lambda_{i})=l_{(i)}\left(1-\lambda_{i}/\overline{\lambda}_{i}\right)$, $\forall i\in[N]$. To justify the robustness of our results, we also test an alternative exponential demand model in Appendix 8.1.
@@
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1036 905 970 674 891 887 666 666 1151 1167 889 598
PG ($\delta_{\text{PG}}=10$) 333 263 281 287 273 292 263 300 608 681 1115 435
PG ($\delta_{\text{PG}}=100$) 371 362 380 400 342 347 358 377 1048 778 723 414
MM 59 59 58 59 66 70 77 74 121 119 187 220
Num. Iterations PG ($\delta_{\text{PG}}=1$) 86.3 76.7 82.7 57.7 76.7 76.0 57.3 57.0 98.3 98.3 76.0 51.0
PG ($\delta_{\text{PG}}=10$) 27.3 22.3 24.0 24.3 23.3 25.0 22.7 25.7 52.0 57.7 95.0 37.0
PG ($\delta_{\text{PG}}=100$) 31.0 30.7 32.3 34.0 29.3 29.7 31.0 32.3 89.3 66.3 61.7 35.3
MM 5.0 5.0 5.0 5.0 5.7 6.0 6.7 6.3 10.3 10.0 16.0 18.7
Objective Function PG ($\delta_{\text{PG}}=1$) 64.05 58.93 45.88 37.99 39.55 33.88 20.67 13.83 16.70 13.22 4.77 0.92
PG ($\delta_{\text{PG}}=10$) 64.26 59.13 46.00 38.11 39.62 33.98 20.81 14.07 12.46 10.28 4.50 1.57
PG ($\delta_{\text{PG}}=100$) 64.26 59.13 46.00 38.10 39.65 33.97 20.85 14.07 13.23 10.50 4.28 1.59
MM 64.26 59.13 46.00 38.11 39.65 33.98 20.85 14.07 19.66 14.42 4.95 1.69
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1200 1200 869 917 1038 720 1000 993 1200 1200 1004 1172
PG ($\delta_{\text{PG}}=10$) 722 770 746 775 701 702 698 727 1185 992 1063 1048
PG ($\delta_{\text{PG}}=100$) 899 891 836 868 873 866 900 980 1203 969 1200 1128
MM 132 133 130 134 159 177 176 159 378 433 541 314
Num. Iterations PG ($\delta_{\text{PG}}=1$) 45.7 45.7 33.3 34.7 39.0 27.3 38.0 38.0 46.0 45.3 38.3 44.0
PG ($\delta_{\text{PG}}=10$) 27.3 29.0 28.7 29.3 26.0 26.7 26.3 27.3 45.3 38.0 40.3 39.7
PG ($\delta_{\text{PG}}=100$) 34.0 33.7 32.0 33.0 33.0 33.0 34.0 36.7 46.0 37.0 45.3 43.0
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.7 6.0 14.3 16.3 20.7 11.7
Objective Function PG ($\delta_{\text{PG}}=1$) 61.95 56.48 44.57 37.76 36.61 30.70 19.47 13.63 15.01 11.46 4.37 1.52
PG ($\delta_{\text{PG}}=10$) 62.42 56.95 44.76 37.96 36.85 31.08 19.72 13.89 14.01 10.36 4.36 1.09
PG ($\delta_{\text{PG}}=100$) 62.42 56.95 44.65 37.96 36.85 31.08 19.72 13.90 13.48 10.31 4.28 1.10
MM 62.42 56.95 44.76 37.96 36.85 31.08 19.72 13.90 16.75 12.13 4.75 1.74
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1132 1128 838 1156 1148 1131 1162 1146 1200 1134 1099 1133
PG ($\delta_{\text{PG}}=10$) 1200 1200 1200 1200 1200 1200 1200 1200 1200 1179 1200 1200
@@
MM 232 234 232 236 285 311 297 286 666 642 661 499
Num. Iterations PG ($\delta_{\text{PG}}=1$) 24.3 24.3 18.0 24.7 24.3 24.3 25.0 24.3 25.3 24.3 23.7 24.0
PG ($\delta_{\text{PG}}=10$) 26.0 26.0 26.0 26.0 25.7 26.0 26.0 25.7 25.3 25.3 26.0 26.0
PG ($\delta_{\text{PG}}=100$) 25.7 26.0 26.0 26.0 25.7 26.0 26.0 25.7 25.7 25.7 26.0 26.0
MM 5.0 5.0 5.0 5.0 6.0 6.7 6.3 6.0 14.0 13.7 14.0 10.3
Objective Function PG ($\delta_{\text{PG}}=1$) 60.25 54.99 18.79 37.66 34.62 29.32 19.01 13.42 14.15 10.87 4.06 1.19
PG ($\delta_{\text{PG}}=10$) 60.75 55.47 44.22 37.81 34.81 29.58 19.27 13.68 13.32 10.26 4.14 1.33
PG ($\delta_{\text{PG}}=100$) 56.67 54.40 43.22 36.91 33.27 29.14 18.37 11.62 10.20 7.16 2.57 -6.12
MM 60.77 55.51 44.27 37.86 34.88 29.59 19.29 13.71 15.28 11.37 4.58 1.66
Table 3 presents the results of the first set of experiments, where we assume $\theta_{1}=\cdots=\theta_{N}=\theta\in\{1/5,1/3,1,2\}$ per minute. Accordingly, the mean sojourn times for riders are $\{5,3,1,0.5\}$ minutes, respectively. The results demonstrate that the MM algorithm substantially reduces both the running time and the number of iterations compared to the PG method. Compared to PG with three different step sizes ($\delta_{\text{PG}}=1,10,100$), the MM algorithm reduces average running time by 76.3%, 70.9%, and 73.0%, respectively, and reduces the average number of iterations by 81.6%, 73.0%, and 75.7%. Since the running time is capped at 20 minutes, the actual time savings may be even greater. Moreover, within the maximum allowed running time, the MM algorithm consistently achieves a higher objective value—on average, 4.31%, 2.17%, and 5.00% higher than the PG methods. We also observe that the performance of PG is sensitive to step size selection: although a step size of 10 yields the best average performance, as shown in Table 4, it is not optimal in every instance: A step size of 1 performs better when $c=1.1$. In contrast, MM demonstrates strong robustness and outperforms PG across all cases and step-size rules.
@@
PG ($\delta_{\text{PG}}=1$) 48.28 25.39 8.19
PG ($\delta_{\text{PG}}=10$) 50.65 25.61 7.31
PG ($\delta_{\text{PG}}=100$) 50.06 25.21 6.05
MM 50.67 25.63 9.08
It can also be observed that as the number of types $N$ increases, the running time of both PG and MM increases, primarily because solving $c(\boldsymbol{\lambda})$ involves $N+N^{2}$ variables at each iteration. However, the number of iterations required by the MM algorithm does not scale proportionally with $N$, and in many cases remains nearly constant at a single-digit value. Even when $N=200$, the MM algorithm remains efficient and consistently produces results within a short CPU time.
@@
$N=100$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 585 559 554 562 581 678
PG ($\delta_{\text{PG}}=10$) 321 282 302 286 831 662
PG ($\delta_{\text{PG}}=100$) 366 394 347 378 1120 709
MM 58 58 70 82 138 183
Num. Iterations PG ($\delta_{\text{PG}}=1$) 50.7 48.7 47.7 48.3 50.3 58.7
PG ($\delta_{\text{PG}}=10$) 27.7 24.3 25.7 24.3 72.0 57.7
PG ($\delta_{\text{PG}}=100$) 31.7 34.0 29.3 32.0 97.7 61.3
MM 5.0 5.0 6.0 7.0 12.0 15.7
Objective Function PG ($\delta_{\text{PG}}=1$) 61.49 41.58 36.64 16.99 15.08 2.97
PG ($\delta_{\text{PG}}=10$) 61.66 41.73 36.79 17.11 12.46 2.62
PG ($\delta_{\text{PG}}=100$) 61.66 41.70 36.79 17.15 11.74 2.02
MM 61.66 41.73 36.79 17.15 16.96 3.09
$N=150$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 899 974 815 842 981 898
PG ($\delta_{\text{PG}}=10$) 764 819 691 659 1051 836
PG ($\delta_{\text{PG}}=100$) 856 857 874 893 1200 1125
MM 132 130 155 166 414 305
Num. Iterations PG ($\delta_{\text{PG}}=1$) 34.0 37.3 31.7 32.7 37.7 34.3
PG ($\delta_{\text{PG}}=10$) 28.7 31.0 26.3 25.3 41.0 32.0
PG ($\delta_{\text{PG}}=100$) 32.7 32.7 33.0 34.3 47.0 42.3
MM 5.0 5.0 6.0 6.3 15.7 11.7
Objective Function PG ($\delta_{\text{PG}}=1$) 59.40 40.88 33.64 16.35 13.38 2.73
PG ($\delta_{\text{PG}}=10$) 59.66 41.15 33.93 16.49 12.41 2.84
PG ($\delta_{\text{PG}}=100$) 59.66 41.16 33.93 16.66 10.97 2.94
MM 59.66 41.16 33.93 16.68 14.36 3.08
$N=200$ Running Time (sec) PG ($\delta_{\text{PG}}=1$) 1200 1031 1137 1148 1200 1197
PG ($\delta_{\text{PG}}=10$) 1200 1200 1200 1200 1200 1200
@@
MM 235 237 278 280 722 479
Num. Iterations PG ($\delta_{\text{PG}}=1$) 26.3 22.0 25.0 24.7 26.0 25.3
PG ($\delta_{\text{PG}}=10$) 26.3 25.7 26.3 26.0 26.7 25.3
PG ($\delta_{\text{PG}}=100$) 26.7 25.3 26.3 26.0 26.0 25.3
MM 5.0 5.0 6.0 6.0 15.3 10.0
Objective Function PG ($\delta_{\text{PG}}=1$) 57.51 40.15 31.91 16.10 12.63 2.60
PG ($\delta_{\text{PG}}=10$) 58.08 40.90 32.19 16.40 11.70 2.32
PG ($\delta_{\text{PG}}=100$) 55.01 36.09 31.25 12.72 4.90 -0.54
MM 58.13 40.96 32.20 16.42 13.35 2.93
Table 5 presents the experimental results for the setting where $\theta_{i}$ varies across agent types. For type-$i$ agents, $\theta_{i}$ is randomly drawn from a uniform distribution $U[\underline{\theta},\overline{\theta}]$, with two parameter choices: $(\underline{\theta},\overline{\theta})\in\{(1/5,1/3),(1,2)\}$. The results exhibit a similar trend to that observed in Table 1. Compared to three PG variants with $\delta_{\text{PG}}=1,10,100$, the MM algorithm reduces running time by 74.0%, 72.0% and 74.7%; decreases the number of iterations by 77.7%, 74.2% and 77.7%; and improves the objective value by 1.63%, 1.95% and 7.23% within the 20-minute maximum runtime, respectively. The performance of PG is still sensitive to the step size: $\delta_{\text{PG}}=10$ performs better when $c\leq 0.9$, whereas $\delta_{\text{PG}}=1$ is preferable when $c=1.1$.
We also conduct simulation studies by implementing the optimal prices given by our MM algorithm and comparing with the baseline pricing scheme when agents’ finite patience is not incorporated (i.e., when (2c) is relaxed, as in the LP formulation of Eom and Toriello 2024). We find that our pricing results consistently outperform the baseline. For details of the simulation, please refer to Appendix 8.2.